# The Dynamic Net Architecture: Learning Robust and Holistic Visual Representations Through Self-Organizing Networks

### v1

@@ Abstract
We present a novel intelligent-system architecture called “Dynamic Net Architecture” (DNA) that relies on recurrence-stabilized networks and discuss it in application to vision. Our architecture models a (cerebral cortical) area wherein elementary feature neurons encode details of visual structures, and coherent nets of such neurons model holistic object structures. By interpreting smaller or larger coherent pieces of an area network as complex features, our model encodes hierarchical feature representations essentially different than artificial neural networks (ANNs).
DNA models operate on a dynamic connectionism principle, wherein neural activations stemming from initial afferent signals undergo stabilization through a self-organizing mechanism facilitated by Hebbian plasticity alongside periodically tightening inhibition. In contrast to ANNs, which rely on feed-forward connections and backpropagation of error, we posit that this processing paradigm leads to highly robust representations, as by employing dynamic lateral connections, irrelevant details in neural activations are filtered out, freeing further processing steps from distracting noise and premature decisions.
We empirically demonstrate the viability of the DNA by composing line fragments into longer lines and show that the construction of nets representing lines remains robust even with the introduction of up to $59\%$ noise at each spatial location. Furthermore, we demonstrate the model’s capability to reconstruct anticipated features from partially obscured inputs and that it can generalize to patterns not observed during training. In this work, we limit the DNA to one cortical area and focus on its internals while providing insights into a standalone area’s strengths and shortcomings. Additionally, we provide an outlook on how future work can implement invariant object recognition by combining multiple areas.
Keywords net fragments  $\cdot$ neural code  $\cdot$ early commitment  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neuronal networks  $\cdot$ neuroscience  $\cdot$ machine learning
@@ 1 Introduction
We present the first implementation of a novel intelligent-system architecture called “Dynamic Net Architecture” (DNA). We motivate it here by contrasting it to the architectures of artificial neural networks (ANNs) that play an essential role in analyzing high-dimensional data (LeCun et al., 2015; Schmidhuber, 2015; Stadelmann et al., 2018). An ANN is a static, deeply nested function optimized with respect to a static cost function. An input to an ANN is mapped to an output by a succession of nested functions arranged in layers. Consequently, representations have the form of a hierarchy that integrates simpler low-level features (outputs of inner functions) to more complex high-level features (outputs of outer functions). The representations computed by the inner functions are independent of the outputs of outer functions. We posit that this is one of the leading causes of the lack of robustness witnessed in ANNs (Windrim et al., 2016; Li et al., 2019; Wang et al., 2019; Rusak et al., 2020; Drenkow et al., 2022) since the decisions of inner functions cannot be adjusted by plausibility checks based on outer functions. This assertion finds support in the formal explanation provided by Goodfellow et al. (2015), who argue that adversarial attacks rely on perturbations below the threshold of visibility when these are summed over a sufficiently high input dimensionality (cf. Ilyas et al. 2019). Following Marr (2010), we call this phenomenon of lacking robustness due to sequential information processing “the fallacy of early commitment,” i.e., the inability to factor the outputs of the outer functions (high-level features) into the computation of the inner functions (low-level features), so that low-level decisions are in danger of misleading high-level decisions.
With the proposed DNA, we overcome this limitation by dynamically integrating local and global features. The DNA is formulated in terms of cortical areas (two-dimensional sheets of neurons) and projections between them. Connectivity within areas is dominated by short-range lateral (“recurrent”) connections, different areas are connected bidirectionally. The neurons of a given area are elementary feature detectors. Afferent connections from a given input pattern activate a fraction of all neurons in the area. Cyclically rising inhibition then silences the majority of the initially activated neurons, leaving only those active that are supporting each other with sufficient strength through excitatory lateral connections. This mutual support can only be realized within specifically structured consistent networks, “nets” (von der Malsburg, 2018) and have to be generated by statistical learning. Nets can span the entire visual field and connect feature neurons that are spatially distant through chains of locally connected intermediary neurons. Nets are flexibly composed of “net fragments” that correspond to often-occurring feature constellations. Each position of an area contains many fragments (analogous to the code-book of image compression schemes, Taubman and Marcellin (2002)) from which the input can choose. Unlike the pieces of a jigsaw puzzle, each net fragment fits to a choice of possible neighboring fragments (and each neuron can be part of several net fragments), so that a virtually infinite number of nets can be formed from a finite set of neurons and net fragments.
According to this principle, a two-phase process admits first a transient initial volley of activity, while in a second phase, most of this activity is silenced so that only mutually supporting neurons stay active. What is responding to the input of a cortical area are thus not individual neurons but rather organized nets. From Gestalt psychology (Wertheimer et al., 1938; Köhler, 1992), we know the visual system can rapidly discern holistic structures, that is, recognize global patterns as arrays of local features that conform consistently to certain “local Gestalt rules” Köhler (1992): from all the local features present in the input, only those are selected by the visual system that fit consistently into a coherent global pattern (coherent in the sense of being composed throughout of overlapping fragments). Thus, local decisions are conditioned on global patterns, whereas global patterns need for their constitution local features. We offer our nets of lateral connections as the mechanism that mediates between local and global levels: local Gestalt rules are implemented by net fragments, and as long as an input pattern can be covered throughout by fragments, the response it elicits is a coherent net.
The two-phase process described is able to remove noise and filter out irrelevant details. It thus avoids the fallacy of early commitment by conditioning neuron activation on integrability into a globally consistent net. An additional effect is that neurons that aren’t receiving afferent input but have sufficient lateral support might get activated, thus complementing figures or textures and allowing the interpretation of occluded objects.
The connectivity necessary for stable second-phase activity is very specific and needs to be generated by learning. The goal is to find and stabilize those patterns of afferent activation that stand out statistically. This is possible with the help of Hebbian plasticity (Hebb, 1949) that strengthens short-range excitatory connections between neurons often co-activated by the input. In the initial stages of learning, before the establishment of supporting connections, learning has to rely on phase one activity, that is, on the actual statistics of input activation. In the course of learning, however, plasticity has to be biased more and more in favor of those neurons that stand out by lateral support, neurons whose activity survives into phase two.
How do our cortical areas compare to the layers of ANNs? Whereas the latter are mainly there to accommodate a hierarchy of feature detectors of higher and higher complexity, already a single cortical area provides for such hierarchy in the form of nested net fragments of growing size (von der Malsburg et al., 2022b), a hierarchy that extends up to input-activated nets representing entire objects. One role of higher cortical areas is to permit the construction of representations invariant to transformations affecting lower areas (translation, rotation, scaling, etc.), while the connection between variant and invariant representations is established by structured dynamic sets of connections (see below).
Areas are dominated by short-range lateral connections, which abound in the cortex Stettler et al. (2002); Suzuki et al. (2023) but are ignored in multi-layer perceptrons (MLPs) (Prince, 2023) and transformer architectures (Vaswani et al., 2017). Lateral, recurrent connections play important roles both for function and for learning. As for function, they are the basis for the establishment of nets as high-level (e.g., object-global) representations by a collective computational mechanism Forrest (1990), conditioning global and local decisions on each other, as noted above. Such mediation between levels is not possible in ANNs with their purely feed-forward flow of information (which is unavoidable due to their learning principle of backpropagation of error (Rosenblatt, 1962; Linnainmaa, 1970; Rumelhart and McClelland, 1987)). Regarding learning, the restriction to connectivity patterns that are self-consistent (connectivity structure supporting activity states that stabilize this same connectivity structure) reduces the search space of learning by many orders of magnitude and has the potential to produce connectivity patterns representing novel structures never seen as examples. Moreover, the stabilization of connection weights based on local rules (rather than backpropagation through many layers) decisively relaxes precision requirements put on them.
In this work, we present the first implementation of a cortical area structured by fragments and apply it to a simple vision task, showcasing its ability to be robust against noise and perform figure completion. We believe that this conceptual work lays the foundation for a novel learning system that can fulfill complex tasks such as invariant object recognition when scaled up to multiple areas in subsequent research (see Section 2.3). Specifically, we make the following contributions:
  • We demonstrate the viability of net fragments as neural code, highlighting their role in enhancing the robustness of computational frameworks.
  • We provide a summary of the relevant neuroscientific hypotheses of net fragments and dynamic linking.
@@
  • We translate these hypotheses into a computational framework compatible with recent machine learning advances.
  • We discuss the introduced concepts and describe how extending it with multiple areas can foster invariant object detection.
2 Related Work
As outlined above, deep neural networks exhibit a deficiency in robustness. In the scientific literature, robustness is typically defined as immunity to noise (additive noise, or subtractive noise in the form of missing information; see Cao et al. (2021); Neururer et al. (2024)) and adversarial attacks (Drenkow et al., 2022), as well as the capability of dealing with ambiguous inputs (Simmler et al., 2021) and generalization within/across domains (e.g., over image translation, rotation, changing lighting, etc.).
This work focuses on analyzing net fragments that are formed in each DNA-based area concerning their robustness to small perturbations. Related work pertaining to this aspect is summarized in Section 2.1, followed by a review of representation learning. The subsequent Section 2.3 examines systems utilizing correspondence-based mapping, which can cope well with ambiguous inputs and exhibit good generalization capabilities. These systems closely align conceptually with the DNA proposed in this work, making them highly relevant for future research, especially for scaling to a multi-area DNA.
2.1 Robustness to Noise
Adversarial attacks exploit deep networks’ vulnerability to noise by calculating slight perturbations capable of altering the model’s predictions for specific samples (Szegedy et al., 2014; Goodfellow et al., 2015; Papernot et al., 2016a; Carlini and Wagner, ; Deriu et al., 2022) or perturbations applicable across the entire dataset (Moosavi-Dezfooli et al., 2017). Adversarial attacks can also be devised without direct access to the model and its parameters, for example, by analyzing a model’s decision boundary (Brendel et al., 2018; Engstrom et al., 2019) or training an auxiliary model (Chen et al., 2017; Cisse et al., 2017; Sarkar et al., 2017).
Current approaches aim to counter these attacks by modifying the network’s architecture to increase robustness (Papernot et al., 2016b; Gao et al., 2017; Ros and Doshi-Velez, 2018), including adversarial examples to the training data (Moosavi-Dezfooli et al., 2016; Goodfellow et al., 2015), transform the inputs before being processed (Dziugaite et al., 2016; Bhagoji et al., 2018), or adding external models (Meng and Chen, 2017; Xu et al., 2018) or attack detectors (Amirian et al., 2018).
The lack of robustness also manifests itself in poor generalization between domains, where a change in the statistical distribution of data prevents the applicability of a model trained on a source domain to a slightly different target domain (Tuggener et al., 2022; Yan et al., 2024). Addressing such domain shifts typically involves domain alignment (Erfani et al., 2016; Ghifary et al., 2017; Rahman et al., 2020; Sager et al., 2022), meta-learning (Dou et al., 2019; Zhou et al., 2020; Du et al., 2020), or data augmentation (Carlucci et al., 2019; Zhang et al., 2020; Tuggener et al., 2024).
Despite notable advances in improving robustness and preventing adversarial attacks, existing methods primarily address symptoms by reducing the number of parameters (Papernot et al., 2016b), removing specific features (Gao et al., 2017), regularizing gradients (Ros and Doshi-Velez, 2018), augment or extend the training data (Moosavi-Dezfooli et al., 2016; Dziugaite et al., 2016; Tuggener et al., 2024), align domains (Erfani et al., 2016), or meta-learn features (Dou et al., 2019). In contrast, we here attempt to fundamentally rethink the principles underlying learning, network architecture, and information flow, potentially leading to more robust representations. In our experiments (see Section 6), we substantiate the efficacy of the proposed net fragments in effectively mitigating random Gaussian noise and dealing with occluded objects (a type of subtractive noise).
2.2 Representation Learning
Other research relevant to this work employs other design principles to enhance robustness. A crucial principle involves sparse neural activation, which enhances robustness relative to dense representations, as shown for deep neural networks (Guo et al., 2018; Liao et al., 2022; Timpl et al., 2022; Prince, 2023). This is particularly true for binary sparse distributions (Ahmad and Scheinkman, 2019). Element-wise comparison of two vectors, each with only a few activated neurons, proves to be highly robust. Given sufficiently large vector dimensionality, accidental similarity across different binary sparse representations is highly improbable (Ahmad and Hawkins, 2015).
Also in this work, we exploit binary sparse representations to bolster model robustness. In our system, sparsity is not imposed by specific constraints but by design along established principles, including lateral connections (Kothari and Agyepong, 1996; Corchado et al., 2003) and self-organization through local Hebbian updates (Hebb, 1949), coupled with inhibitory mechanisms for activity regulation (Abbott and Nelson, 2000; Luz and Shamir, 2012).
2.3 Correspondence-Based Mapping
A highly robust vision framework for object detection is based on correspondence mapping between pixel patterns within an image and a corresponding prototype (Wolfrum et al., 2008). Here, “corresponding” refers to neurons relating to the same point on the object’s surface. A pivotal concept in this framework are maps implemented as sets of projection fibers (connections that dynamically turn on and off) known as shifter circuits (Anderson and van Essen, 1987; Olshausen et al., 1993) that are composed of maplets (Zhu and von der Malsburg, 2004). These maps are activated when point-to-point correlations are detected between coherent nets in two cortical areas. The composite nets in the two cortical areas explicitly represent spatial relations, and a map between them can only be formed if they agree on this spatial structure. Once a forward mapping to a coherent model of a visual pattern is established, backward mapping to the primary visual domain can stabilize the modeled structure there by complementing missing elements or by lacking to support visual noise elements.
So far, correspondence-based mapping has been applied only to human face detection and identification (Lades et al., 1993; Wolfrum et al., 2008; Fernandes and von der Malsburg, 2015), in which image and object model match with little deformation. For general classes of visual structure, mapping has to tolerate deformation and must be based on the matching of robust feature constellations (Biederman and Kalocsai, 1997). Net fragments can be seen as an attempt to model these (von der Malsburg, 2014, 2018; von der Malsburg et al., 2022a). So far, however, net fragments have only been proposed as a theoretical concept and have not been modeled in any concrete fashion.
In this work our focus is, however, not on correspondence-based mapping, but on the learning of fragment-based net formation, hoping to thereby contribute to the establishment of this neural code as a sound conceptual framework translating concepts of biological vision to applications in next-generation machine learning systems.
3 Net Fragments for Visual Perception
In the following, we restrict the discussion of DNA to the visual system. Light fields captured by the human eyes are translated to retinal images that appear as two-dimensional arrays of activated neurons in the primary visual cortex (Grill-Spector and Malach, 2004). There, neurons are sensitive to local texture elements within their receptive fields (Grill-Spector and Malach, 2004) and interact through short-range lateral connections (Gilbert et al., 1990). Visual patterns that appear with significant frequency within an image patch lead, through Hebbian plasticity, to the formation of net fragments, as visualized in Figure 1.
Figure 1: Net fragments in a nutshell: The eyes observe a visual scene and activate neurons in the primary visual cortex. Neurons that support each other inside net fragments remain active, while neurons initially activated by afferent excitation but lacking such support are silenced by inhibition.
Among afferent input-activated neurons, only a few are part of net fragments so as to support each other’s activity by exchanging spikes along excitatory lateral connections (Stettler et al., 2002). Within milliseconds, neurons that have sufficient lateral support are singled out by neural inhibition (Vogels et al., 2011) from all others that also receive afferent input but lack lateral support.
All statistically significant visual textures within image patches with the radius of the range of lateral connections are represented by local connectivity patterns (such as in image compression algorithms (Wallace, 1991) by codebook vectors). To account for the homogeneity of image structure, it can be assumed that the connectivity has translational symmetry (“convolutional” structure) supporting equivariance, that is, the property that image shift gets reflected in activity shift.
A given feature can be part of a large number of patterns. To avoid interference between these patterns, each feature (defined in the primary visual area by a particular receptive field shape in a particular position) is represented by a number of “alternative neurons,” which are free to carry different connections  [^1] Alternative neurons may be seen as analogous to transformer’s Vaswani et al. (2017) token embedding vector components, which support the different connections between tokens represented by the query-key mechanism.. In this way, features can vary their connections to other features by activating different alternative neurons: they are connected by “dynamic links.” For more details on alternative neurons, please refer to Appendix A.
To project net fragments and coherent nets out from the mass of all neurons that are individually activated by the sensory input needs dynamic processing: After neurons are activated by afferent input, they are subjected to progressively increasing inhibition, which eventually silences all those of them that don’t have sufficient lateral support. As many fragments can survive inhibition only if supported by overlapping other fragments, inhibition-resistant responses constitute nets that are coherently composed of overlapping fragments.
These fragment-composed nets differ fundamentally from associative memories as formalized, for instance, by Hopfield (1982); Cohen and Grossberg (1983). In these, entire neural patterns, “assemblies,” are stored (in a single step) and can be retrieved only as a rigid whole, while nets are flexibly composed of overlapping fragments (which have been learned in a statistical process) and are flexibly assembled to model the input. We discuss this distinction in more detail in Appendix A.

### v2

# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

@@ Abstract
Nets, cooperative networks of neurons, have been proposed as format for the representation of sensory signals, as physical implementation of the Gestalt phenomenon and as solution to the neural binding problem, while the direct interaction between nets by structure-sensitive matching has been proposed as basis for object-global operations such as object detection. The nets are flexibly composed of overlapping net fragments, which are learned from statistical regularities of sensory input. We here present the cooperative network architecture (CNA), a concrete model that learns such net structure to represent input patterns and deals robustly with noise, deformation, and out-of-distribution data, thus laying the groundwork for a novel neural architecture.
Keywords net fragments  $\cdot$ neural code  $\cdot$ pattern recognition  $\cdot$ computer vision  $\cdot$ neural networks  $\cdot$ cooperativity  $\cdot$ neuroscience  $\cdot$ machine learning
@@ 1 Introduction
Figure 1: Representation of Objects by Nets. Objects are captured in their entirety by coherent nets that are composed of overlapping net fragments (schematically symbolized by white circles in the right panel). After training on natural images, each patch of visual space contains a complement of net fragments that represent textures that have been encountered with statistical significance. Net fragments overlap neuron-wise, and those activated by an object coalesce into a coherent net. The conundrum (stated in Olshausen and Field (2005)) that object contours can often not be found with the help of edge detectors due to lack of gray-level contrast (as in places inside the square in the left panel) may find its solution in that contours are defined as borders of the nets covering objects (or covering the background). Figure adapted from Olshausen and Field (2005), with permission.
Artificial intelligence (AI) systems have demonstrated remarkable capabilities in processing diverse types of data, including images, text, and speech (LeCun et al., 2015; Simmler et al., 2021). While many recent AI systems predominantly favor neural representations over symbolic ones, we follow a proposal by von der Malsburg et al. (2022) and present the cooperative network architecture (CNA), an architecture that uses structured networks of mutually (‘recurrently’) connected neurons as representation of sensory signals. Using structured networks to encode information brings a diverse set of advantages, including high robustness, ability of figure reconstruction, and representing of out-of-domain data, laying the foundations to build novel object recognition frameworks (see Section 6).
The networks in the CNA are called nets and are dynamically assembled from overlapping net fragments, where net fragments are connectivity patterns that are learned by synaptic plasticity based on the statistics of sensory input signals. A major argument in favor of this data architecture is the possibility (as described repeatedly by Hinton (1981); von der Malsburg (1981); Anderson and Van Essen (1987); Kree and Zippelius (1988); Olshausen et al. (1995); Wiskott et al. (1997); Wolfrum et al. (2008); Memisevic and Hinton (2010); Arathorn et al. (2013)) of solving the invariant object recognition problem by structure-preserving mapping between nets representing variant images and invariant models of objects. The representation of image structure by recurrently connected nets formed between elementary feature detector neurons is a fundamental deviation from the current artificial neural networks (ANNs) paradigm, according to which image structure is represented by sets of complex-feature neurons activated by purely feedforward connections (see Section 2.3 for a comparison of the two approaches).
In the cited models of invariant object recognition, the formation of object-representing networks within the image domain (the primary cortical area where sensory signals arrive – here called ‘area’) has been sidestepped by letting spatial neighborhood stand in for connectedness. Neighborhood, however, is too weak a condition to qualify neurons as having responded to the same object (which is, for instance, apparent when looking at an object through foliage). Following von der Malsburg et al. (2022), we rather assume that for sets of neurons to be admitted into a net, they have to have been co-activated frequently in the past by structured visual input and consequently have been tied together by Hebbian synaptic plasticity (Hebb, 1949) into net fragments. Net fragments play a role somewhat related to the code-book vectors of some image compression schemes (Taubman and Marcellin, 2002) in representing statistically dominant texture patterns within any patch of the visual field; the formation of net fragments is somewhat related to, though fundamentally different from, associative memory (Hopfield, 1982), see Section 2.1.
CNA maintains that when a region of the area is hit by signals caused by an object, it transiently activates a subset of all the neurons in the area. The majority of those neurons are not part of a co-activated net fragment and have to be silenced. This is achieved with the help of some periodically rising inhibition, which silences those neurons and leaves active the small subset of all transiently activated neurons that are supporting each other by excitatory connections. Due to neuron-wise overlap, net fragments merge into a coherent net, which can span the entire region covered by an object’s image and thus connect feature neurons that are spatially distant through chains of locally connected intermediary neurons, see Figure 1. It is this connectedness that unites neurons activated by an object in forming a structured net that separates those neurons from the background and from occluding elements of other objects, thus solving the segmentation problem (von der Malsburg, 2024) and tying them into a network amenable to be mapped to an invariant model network in another cortical area (see the outlook in Section 6).
This study introduces the cooperative network architecture (CNA), representing the first concrete implementation of some fundamental components proposed in von der Malsburg et al. (2022). We focus on how to leverage networks as neural code and, specifically, discuss the
@@
  • establishment of a computational framework comparable to recent machine learning advances,
  • learning of net fragments,
  • demonstration of the activation of object-spanning networks (nets),
  • demonstration of filter properties afforded by nets,
  • discussion of nets in contrast to established ideas in neural computation and machine learning,
while relating to the significance of nets for invariant object recognition in multi-area settings only as outlook in the discussion section (i.e. we focus here on a single area while object recognition based on multiple areas is proposed for future research). We demonstrate that net fragments can be learned without relying on a teaching signal, whether supervised or unsupervised. This neural code exhibits robust capabilities for handling out-of-distribution data and effectively represents objects composed of fundamental features encountered during training—a task at which autoencoders (Hinton and Salakhutdinov, 2006) fail.
2 Relation to Alternative Concepts
The model we describe below is the first to demonstrate in concrete form the generation of structured nets that form a neural code to represent visual structure. To grasp the significance of the approach, it is important to compare it to alternative approaches.
2.1 Associative Memory
Associative memory (AM), as formulated by Amari (1972) or Hopfield (1982), shares with the present study the aim of retrieving previously encountered structure, and in both, the structure is recorded with the help of Hebbian synaptic plasticity. Whereas AM deals with the storage and retrieval of neural firing patterns, our purpose is to activate networks, that is, connectivity patterns that are structured by previously recorded connectivity pattern fragments. Admittedly, in both cases, the (re)construction process results in the activation of subsets of all neurons of the area, but whereas in AM, the connectivity of the set of active neurons is all-to-all, in our case, connectivity is restricted to within the net fragments that have previously been recorded, which in our visual application are restricted to short-range (local) connections and express two-dimensional topological structure. It is this structured connectivity that serves as basis for the structure-sensitive matching mentioned in the introduction. The monolithic all-to-all connectivity of the AM, by distinction, does not afford such structure-sensitive matching (i.e. instead of representing each structure as monolithic whole, the CNA learns to decompose them into local sub-patterns (net fragments), which can then be dynamically composed into nets to represent structures).
This functional difference is conditioned on a different mode of learning. Whereas in AM, a neural pattern is stored in a single event by strengthening the connections between any pair of currently active neurons, our approach distinguishes significant from insignificant connections, significant being those that are strengthened again and again, while insignificant are those that are strengthened only occasionally. This learning is driven by exposure to a large number of patterns, none of which may ever repeat, but which contain small repeating subpatterns. It is these that generate net fragments. In the visual case, it is local texture patches that play this role of repeating sub-patterns.
A basic limitation of AM is that unless the patterns to be stored are statistically independent (orthogonal), that is, have small neuron-wise overlap, they interfere with each other by the spurious activation of neurons. The problem can be reduced (as proposed by Tsodyks and Feigel’man (2007) or Palm (2013)) if patterns are sparse, that is, activate only a small fraction of all neurons. The statistical mode of pattern storage proposed here opens a more fundamental solution to the problem of pattern overlap. We propose that each input feature (specified by feature type and position) is represented not by one neuron but by a number of ‘alternative neurons’, neurons that have the exact same receptive field (for examples, see Cossell et al. (2015)), and thus are activated by the same afferents  [^1] There is ample space for alternative neurons, as there are 100 times more neurons in primary visual cortex than output neurons in the retina (Leuba and Kraftsik, 1994).. While these neurons have the same afferent input, they can exhibit different excitatory output. Given this expansion, different net fragments with feature-wise overlap can reduce neuron-wise overlap by making use of different alternative neurons.
AM and our approach differ in yet another aspect, the mode of retrieval (or, in our case, construction) of patterns. In the version of Hopfield (1982) of AM, this is realized by minimizing an energy function by sequential switching of neurons. In our approach, winning patterns are projected out from the set of all neurons activated by the input and by rising inhibition that silences all those alternative neurons that do not have sufficient recurrent excitatory support, no neuron-sequential process being necessary.
2.2 Boltzmann Machine
Some important aspects of the system illustrated by our model are highlighted best by comparing it with the Boltzmann Machine (BM) proposed in Hinton and Sejnowski (1983). Both systems share features such as the ability to admitting recurrent connections (as does Mumford (2003)), but they also exhibit significant differences. Nevertheless, as our focus is limited to one area and is not the full architecture as proposed in von der Malsburg et al. (2022), we highlight only select differences here.
The BM makes a clear distinction between visible and hidden (‘latent’) neurons. In the CNA this distinction is softened: it posits neurons that are activated by sensory input, but of which the majority are quickly silenced by inhibition in favor of those neurons that support each other through previously learned connections. A winning alternative neuron can be interpreted as the hypothesis not only that the neuron has detected its trigger pattern in its afferents but also that the neuron has connections to a number of other active neurons and is consistent with them in light of previous experience.
In both BM and our model, the activity of hidden, or in our case alternative, neurons, is controlled by the same type of dynamic equations, with the difference that we do not insist on symmetric connections and therefore do not derive an energy function and consequently cannot prove convergence of the neural dynamics (see, however, the remark made in Hinton and Sejnowski (1983), that given sufficiently redundant support of active neurons, the missing of up to half of the connections due to lack of symmetry does not jeopardize stability of a state). In a single cycle of the periodically rising inhibition, convergence is not an issue, as alternative neurons undergo winner-take-all competition.
2.3 Artificial Neural Networks (ANN)
In the concept pursued here, objects and other structures within the input are represented by nets, sets of neurons that are held together by cooperative excitatory connections and are activated as a whole, the activation of individual neurons being conditioned on the activation of other neurons in the same net. In contrast, a large variety of ANN versions and learning techniques have been proposed, including multilayer perceptrons (MLPs) (Prince, 2023), transformers (Vaswani et al., 2017), autoencoders (Hinton and Salakhutdinov, 2006), or representation learning (Bengio et al., 2014). Common to all of these is that the sensory input is converted through many layers into a representation in the form of a set of neuron states. These neurons can be considered as propositions that indicate the presence of complex features (activity patterns) within the input. This representation of the input is then exploited by evaluation circuits (often in the form of a few fully connected layers of neurons) to achieve functions such as object classification, next-token prediction, or, in autoencoders, reconstruction of the input. The common feature of all ANN systems is that the connections from the input to the representation layer and from there to the output are purely feedforward and are trained by back-propagation of error (Rosenblatt, 1962, p. 292).
The CNA proposes a clear separation between input representation (within an area) and its functional evaluation (between areas, see Section 6). The representation of the input is trained by its statistics and is thus essentially neutral and without specialization to any specific function, giving it foundational character. In ANN systems, the selection of patterns to be extracted from the input is strongly biased to serve the trained function, which may be the reason that a transformer trained on one set of functions cannot continue to be trained on another (Dohare et al., 2024).
ANN’s hierarchy of layers of neurons responding to features of intermediate complexity is replaced in the CNA by a single layer, in which net fragments of different sizes represent a hierarchy of features of varying complexity (and inter-mesh to form coherent nets standing for whole objects, thus solving the binding problem). The purpose of transition to other areas is, in the CNA, not transition to higher complexity but transition to representations that generalize over, e.g., position, size, or orientation of nets in lower areas, as discussed in Section 6.
In ANN systems, the high-level feature neurons that are activated by an input do not express any relations between them. This ‘binding problem’ (von der Malsburg, 1981; Burwick, 2014; Greff et al., 2020)) can give rise to ambiguity, e.g., the hallucination of feature conjunctions (Treisman and Schmidt, 1982). This ambiguity can be reduced by additional neurons whose trigger patterns overlap those of other neurons whose relative position would otherwise be in doubt (autoencoders (Hinton and Salakhutdinov, 2006) show that ambiguity can be suppressed altogether by a sufficiently large number of hidden neurons). In the CNA, the binding problem is solved in a natural way by net fragments that interconnect sets of features that co-occur with statistical significance in the input and that merge together into nets representing coherent chunks within a given input. This simplicity of representation of visual structure in CNA is due to its fundamentally different perspective on neural representation, in that the active state is not only described by the elementary features that are active but also by the connections between them. This is made possible by different alternative neurons that represent the same elementary feature but express different connectivity.